Telescope.org
This website uses cookies. To read more about our use of cookies, click here.
To not be bothered by this message again, click the OK button.
OK

Project News

Submitted by John on Sunday 2 January 2011

All systems working well. We are currently preparing for the Stargazing live programme with Brian Cox on BBC2 8pm on the 3rd 4th and 5th of January 2011

Submitted by John on Sunday 21 February 2010

We are not using this currently we are using the current status to deliver information.

Submitted by John on Monday 24 November 2008

After failure of two critical weather sensors and the network switch we sent out a replacement switch urgently only to have it take two weeks arriving just in time for John Baruch's management committee visit.
Both are now fixed and the system is working.

Submitted by chris on Saturday 1 November 2008

On the day we say goodbye to the old Java applet V1 for prehistoric JVMs, we also say hello to the new web site image processing system to incorporate flat field use. It's currently in beta testing, but is available to all. You will find the new image processing links under the regular links for viewing / downloading telescope images.

With the current (one) flat field that is installed on the web site, it seems to do a reasonable job of smoothing out the vignetting and removing most of the hairline CCD scratch.

Discuss successes/failures in the forum!

Submitted by John on Wednesday 16 July 2008

The broken gearbox has now been replaced and the system is working well. We are fine tuning as we go along to improve performance.

Submitted by wsmarley on Sunday 1 June 2008

We are currently experiencing a problem with the dome mechanism, meaning that the telescope has been unable to operate for a few nights. More information will be posted as we investigate the situation. Check out this forum thread for more details.

Submitted by wsmarley on Friday 30 May 2008

We are currently experiencing a problem with the dome mechanism, meaning that the telescope has been unable to operate for a few nights. More information will be posted as we investigate the situation.

Submitted by chris on Thursday 13 March 2008

Last night marked the first night of normal operation after the recent servicing. Although the weather wasn't great and ruled out most of the night, some images were taken. They show that the C14 telescope is pointing and tracking well.

Submitted by chris on Saturday 23 February 2008

As you may have guessed, the web site has received a
long-overdue style overhaul. The change also brings new
features:

* The return of the forum

* A real-time data screen showing telescope operations -
watch out for this becoming active when the telescope
starts operating again

* Flash videos of the stars time-lapse videos

* A new webcam - Observatory-Cam

* A list of waiting jobs with their scheduling data

Also as we are now using a wiki based article system for
articles and help on the web site, the help and documentation
part of the website should grow in the coming months.

Enjoy!

Submitted by John on Wednesday 30 May 2007

Today we are celebrating the completion of 30,000 observing requests. We are operating effectively with about 1000 requests of which it is possible to schedule 300 or so and we are completing over 100 requests per night.

Submitted by dan on Saturday 17 March 2007

Occasionally (16/17 March 2007 )there are images with a green tinge or more. As far as we can see this is due to a laser guiding system used between the Island of La Palma 140 km away and the site here in Tenerife.

Submitted by James on Friday 15 December 2006

The Royal Astronomical Society has released the following endorsement for robotic telescopes in Great Britain.

'The Royal Astronomical Society endorses robotic telescopes as an exciting way to give students direct and personal experience of science.
The Society commends the Robotic Telescope Project at Bradford University, the Faulkes Telescopes Project at Cardiff University and the National Schools Observatory at Liverpool John Moores University for bringing to classrooms throughout the UK the results from first-class telescopes in good astronomical sites on distant mountain tops and for providing soundly-based scientific projects so that teachers can use
them for a range of educational needs'.

David Elliott
Executive Secretary

Submitted by James on Wednesday 15 November 2006

The Bradford Robotic Telescope project has a new sister site http://schools.telescope.org/. The website is focused entirely on teaching with a simplified easy to use look for pupils and improved interfaces for class creation and pupil management for teachers. The Educational materials are designed to help teachers deliver the English National Curriculum with scintillating lessons which we hope will particularly help those whose astronomy is a little rusty.

Submitted by James on Wednesday 2 August 2006

New links off the BRT home page

First Contact
Not used the telescope before? First Contact is our way of helping you out: a range of observations using pre-set parameters, which makes them as easy to take as possible.

Reviews and Articles
The number of reviews and articles on the telescope are starting to grow and so we thought it would be nice to share some with you.
If you know of any others please let us know.

If you experience any problems with these new pages or find any errors in them please let us know.

Submitted by James on Friday 14 July 2006

Thanks to everyone that took part in the "I see the Moon" poetry competition. We will post the results on the website soon.

Submitted by chris on Saturday 8 July 2006

Reminder: As posted on the status page, the telescope.org web site will be unavailable from Saturday 07:00 BST. It will be restored probably Saturday evening, but if there is a problem it could be as late as Monday Morning.

Submitted by James on Thursday 29 June 2006

There is still time to get your "I see the Moon" poems in by email before the competitions closing date of 30th of June. Click here for more information

Submitted by John on Monday 19 June 2006

It appears as if the Galaxy camera problem is wider. It looks as if the USB hub on the mount has died. We are working urgently to get it replaced. 19/6/2006

Submitted by dan on Friday 16 June 2006

GALAXY CAMERA OFFLINE

We have experienced problems with the operation of the telescope over the last couple of nights. Investigation is pointing towards a problem with the Galaxy Camera. This camera will be taken off line whilst an investigation of the problem and then solution are found.

Submitted by dan on Monday 29 May 2006

Galaxy Camera Fixed
After a session of cleaning the shutters and testing the galaxy camera appears to be happy with short exposures again. (this was done last night)

Const Camera Focus Fixed
We have adjusted the focus on the constellation camera, images should now be in focus again. (unless you use the none filter option)

Dome Webcam Fixed
The dome webcam has been replaced with a low light B&W camera with an auto iris lens. It should be better around dusk and won't have the tracks of the sun burnt into the CCD.

Submitted by dan on Tuesday 23 May 2006

Hardware troubles

We currently have a problem with the focus on Constellation cam. This camera has been taken off line pending repair.

There is also a problem with Galaxy cam preventing short exposures. The shutters are not fully opening with exposures under 50ms. I have added code to prevent the processing of very short exposures on Galaxy until we can repair the camera.

Trip to repair cameras
In the next few days we will be heading to the observatory to repair the faulty cameras.

Submitted by James on Friday 12 May 2006

SETPOINT West Yorkshire and the University of Bradford invite West Yorkshire Primary Schools to take part in the 'I See the Moon' poetry competition. Click here for more information

Submitted by dan on Friday 24 March 2006

Jobs returned with status 'Err: h/w unable.' from the night of 22/23 March

Due to an error with the focuser on the Galaxy camera on the night of 22/03/06 around 280 jobs were returned with the status h/w unable. The problem with the software has been fixed and all of the jobs will be returned to active status and should be processed as normal.

Submitted by John on Monday 27 February 2006

The weather is back to normal February weather and we are observing normally. Most of the winter damage is repaired but not all and the tracking update is not complete. We expect to complete both these over Easter.

Submitted by John on Saturday 28 January 2006

The weather is continuing to be awful. Since the storm at the end of November there have been few dark hours which have not had high winds, cloud or ice. The system is working even though some of the weather sensors and cameras are damaged. Let us hope the weather soon improves.

Submitted by John on Friday 13 January 2006

We have a new home page. It was only when we were told that we had a fantastic inspirational telescope and learning support which was hidden behind a dull home page that we thought we had to do something. I hope the new home page is a better reflection of what is provided on the site.

Submitted by chris on Tuesday 10 January 2006

Apologies for yesterday's web site down time (9th January). The main web server decided to invent a whole load of i/o errors. It staggered on for a while until I shut it down to investigate. After a few hours of futile investigation I started it up again, and it works...

Submitted by John on Wednesday 14 December 2005

Following tropical storm delta the telescope is operating again. The wind sensors cease to work at wind speeds of 60 metres per second (216km/hour)and they recorded gusts of 55 metres per second (198 km/h). A number of the systems have been broken. Two cloud sensors seem to have had wires ripped out of them. Two web cams have moved - Pole star cam and ESO cam - and Dome cam seems to be broken. The major problems seem to have arisen due to one network switch and the Galaxy cam focusser being broken. We can only assume that is the product of the mains power being cycled so many times.

We have been helped by the loan of a network switch from Jodrell Bank, we are still working on the other problems and ordering the spare parts. We will continue to work with Constellation Cam and Cluster Cam for the time being, which continue to give excellent images.

Submitted by John on Wednesday 30 November 2005

The tropical storm Delta went through the Observatory in the evening of the 28th November 2005. Between 6pm and 8pm the wind was gusting to 50metres per second. At that point the link to the site went down. At the time of writing 30th November 11-30am it has not been reinstated. We understand that there is some damage and we have received the following from the Observatory:

This is to inform you that the Observatorio Teide(OT) is OK after the tropical storm Delta.
At present the OT remains closed and there is no external electrical power but all the installations are OK and working with power generators. 11 persons are located (and right!) at the Residencia. Phone lines (even mobile phones) and internet connection are down and the only possibily to connect with the Observatory is using radio stations from La Laguna. Tomorrow morning we will end the emergency situation and we re-open the Observatory.

Submitted by John on Friday 28 October 2005

We have completed the Winter refit for the telescope and upgrade. Hopefully it will survive. Our plan now is to run it at least two nights per week weather permitting which should be a Monday and a Wednesday. We recognise that this is essential for all our school users and will work hard to ensure that schools always get their requests back within a week; that is if the object is above the horizon at night during that period.

Submitted by dan on Monday 3 October 2005

We have been testing new controlserver code in order to bring on line the long anticipated eves dropping system (the core of which we refer to as DataStore) In the process of the update we have uncovered a problem with the camera control classes that has proved elusive in the tracking down and rectification of. As a reaction to this, in order to keep jobs moving through the system we are planning to run an older version of the controlserver software until the problem with the new code can be fixed. This should not cause problems for the user, as the code changes has mainly been to backend systems.

Submitted by chris on Thursday 14 July 2005

The web site has been updated to include some new features and document structure. The new features include the ability to search all the jobs in the database, view any job in the database and submit good images from the applet to a new JPEG image gallery on the web site.
The new document structure of the site gives us room to add help and FAQ pages and other interesting informational content.

Submitted by dan on Wednesday 15 June 2005

We have now added focus offsets for the filters of Galaxy Cam. This will hopefully yield better focussed imaged. The initial testing looks good, so the clear filter on Galaxy Cam is now usable.

Also we have a new pointing model. It is rough at the moment, but should improve over the next few days.

The system is now running jobs again. (Though as we are still on site, the activity may be intermittent.)

Submitted by chris on Wednesday 8 June 2005

Well after a long time with no updates, some of the worst weather on the mountain that there has been for years, two trips to the site to fix broken hardware and a lot of coding, the system is now operating smoothly. Therefore, the website is hereby declared open! No beta password is required to create an account any more.

Don\'t take this to mean that the system is 100% finished and working perfectly - there are still many issues to resolve, features to create, etc. But we are now at the stage where we need more users and more job requests entered into the system to keep it busy.

Also, having just opened the web site to new accounts, we are going to be taking the telescope offline in a few days time for re-alignment, but all jobs will be queued and done with hopefully greater accuracy just a few days later.

Submitted by dan on Wednesday 26 January 2005

We are currently testing new dome control code. This new code will bring a number of enhancements...
1) Greater reliability
2) More efficient operation
3) Per camera offset (images with less dome, more sky)

To test the new code we are placing user jobs on hold. So if you see any of your jobs on hold, you are not being singled out!! it is just testing. :)

The new code should come online any time now, and then the telescope will be back to operating full time.

The new code uses entirely new way of positioning the dome and we expect that there might initially be problems, especially when near the meridianÂ… if this happens to you please let us know.

dan

Submitted by chris on Monday 6 December 2004

The telescope system has not been running for some time due to bad weather, and now it will not run for approximately the next week or so due to major upgrade work.

Submitted by dan on Saturday 4 December 2004

Pole star and Teide star cameras off-line pending major upgrade. (expected in next two days)

Submitted by John on Wednesday 17 November 2004

The Job ID 1187 with Request ID 1456 is a beautiful image of Cassiopeia along with clusters of stars and illuminated clouds in the sky.

Submitted by chris on Thursday 30 September 2004

Good news, we have installed a replacement UPS and power supply for the control server, we have our network connectivity back and everything should be working as before.

Submitted by chris on Friday 24 September 2004

STOP PRESS... At 04:48GMT 22nd September 2004 we lost contact with the telescope site. Further testing shows us that the Control computer power supply probably failed and brought down the UPS support. Our flexibility means that we have some software redundancy but there are some key parts of hardware that need hands on reconfiguration.

In the next few days we are planning more beta testing with local primary schools and it is vital that we get on line again.

Chris and Dan have volunteered to make a special trip out to get everything running again.

Thank goodness it is not Australia. This is the first time a power supply has failed in about 15 computer-years of operation.

John

Submitted by dan on Wednesday 30 June 2004

New good/bad rules for cloud sensors in testing. (visible on the weather reports page)

Submitted by dan on Wednesday 30 June 2004

On Sunday night we returned from a successful trip to the Tenerife observatory.
Main points of work carried out on trip:
Final setup of FLI cameras.*
Long alignment run.
Analemma cam lens change and new solar filter.
New sun rise and sun set light sensors.
Switch to VMware for pointing OS.
New automatic red light in dome.**
New lenses on web cameras.

* Due to the large number of electrical and rf systems in the observatory we had some noise problems with the long cable runs controlling the cameras. With support from FLI with fixed the noise problems, then upgraded to their new ME2 cameras. Noise levels are now very good.

** Red light in dome computer controlled. This allows us to turn the light on whilst not imaging to see the system at work. When/if this system is introduced the pier cam will be illuminated during the slewing of the telescope.


At the end of the visit we conducted testing of the autonomous system. With the new alignment model we have a number of images from this test confirming the pointing model accuracy. Images may be posted here later.

Submitted by dan on Thursday 6 May 2004

There is a new updated weather server in final testing at the moment. This may cause some problems with data logging to the database. It has been tested in the lab, but just a quick warning.

This update basically incorporates a number of changes, the biggest of which is a restructuring of the goodbad system to use threads. This will make the system a lot more efficient. Other changes include modification of the timers code to cope with the new method of system time synchronisation.

Submitted by chris on Tuesday 20 April 2004

As mentioned before, we did not get the C14 aligned due to bad weather in the December trip to Tenerife. We have now completed another trip to Tenerife where the weather conspired against us again but we did get enough alignment and focussing done to use the C14 remotely. So, after a few last pieces of software work now that we have the system ready to go, we should be able to open for beta testing fairly soon!

Submitted by chris on Wednesday 24 December 2003

After another mostly successful trip to Tenerife, we now have a Celestron C14 telescope almost ready to operate, and a constellation camera that will be ready to operate soon. Although the alignment work for the C14 was hampered by bad winter weather we will get that done next time we go over there.

Submitted by chris on Friday 17 October 2003

We've done it!

After some setbacks attempting to control the mount which we have now overcome, we have taken our first job through the system from website to telescope and back to website.

Submitted by chris on Monday 6 October 2003

There has been quite a gap since my last post but now I have good news to report.

During the time since we last visited the observatory on Tenerife we have been gradually ordering more parts for the system, writing software and waiting for the dome firmware. On arrival of the new dome firmware two weeks ago we headed out to Tenerife to get the observatory working. The two most important jobs we had to do out there were successful: the dome software update worked and did what we needed, and we completed a 60 point mapping run for the telescope. This means that in the event of power or computer failure our dome will close (rather important) and the second job means that the mount is aligned with the sky well enough to point the telescope properly. On top of these two critical jobs we also did the following:

Installed a new power supply for the internal weather sensor junction box in the fight against weather data spikes.

Installed a new power supply for the weather sensor heaters - they were drawing too much current from the system and were affecting some new sensors we tried to install a couple of trips ago.

Applied some ferrite cores to weather sensor cables to reduce the spikes.

Using a rabbit board we installed a small embedded system that pings all of our computers on site over the network and reboots them if they do not respond for a given time. We can also log into the reset box and reboot any computer manually. This will allow us to recover the system remotely if a machine should crash. Although this hasn't happened yet in the year we have had our new machines running out there, it might do one day.

We replaced some computer fans that were already making unhealthy noises, painted the rest of the floor, installed a wireless lan access point and took some sample images of the sky with different cameras.

On the webcams front: Before this trip we had 5 cameras, 4 of which entered into a 4 way switch box and then into a computer. The other camera went straight into another computer. This trip we installed the new Teide-Stars-Cam and Analemma-Cam - to accomodate this we also installed another 4 way switch box on the other computer. We now have 7 cameras with one input spare. Finally we fixed the software which does the image processing for Pole-Star-Cam - many stars can now be seen.

Overall a successful trip as we now only have to complete some of the safety software before we can start testing the installation remotely.

Submitted by chris on Monday 28 July 2003

The forums are working again. Specifically, the forums suddenly decided to disallow new signups due to some obscure database error. This is all fixed now. Make yourself an account, be nice!

Submitted by chris on Monday 14 July 2003

Guess what! Yep, it's results from another Tenerife trip time!

Ok, this time we didn't take a long list of small jobs to do, we took a short list of long jobs, and unfortunately, we weren't all that successful with some of them.

First of all, the bits that went ok:
Dan:
- re-installed the dew sensor on the new weather system and located it in the dome.
- fixed a problem with the internal weather junction box that may have fixed the spikes, the only way to tell is to wait and watch the results.
- worked out what was wrong with the external temperature sensor and fixed it.
Chris:
- wrote some test software for the focusser
- tried some filters in roadcam to see if we could see the outline of the sun
- wrote the job de-schedule code
Chris and Dan:
- worked on the main control server and got quite a lot of it glued together and working
- got the Paramount to stop giving us strange errors and failing to do its job
- got the Paramount to work properly through our Moxa box
- did a 25 point alignment run

Now for the bit that went wrong...
We need our dome to close on its own if there is a power cut or a computer crash; the software to do this was being written by the dome automation supplier. We didn't know we would get it in time for this trip so Dan started building a work-around while we were in Tenerife. While we were there though, we got notification that we would receive the dome updates so we turned out attention to other parts of the site. The updates arrived and correctly closed the dome when necessary but unfortunately they had lost the ability to count to 365 degrees per dome rotation, so we had to go back to the old version of the software. Unfortunately by this time we didn't have enough time left to implement the dome close solution.

So, the next trip should be the one-where-we-get-it-working. (again). The dome updates will hopefully be sorted, we will hopefully install a Celestron tube with associated camera and filters, the control server is practically working already, communications from base web site to telescope site is done and basically, everything looks pretty much nearly finished.. Watch this space!

Submitted by chris on Monday 30 June 2003

Results from another Tenerife trip:

We installed:
A pier adaptor
Paramount ME mount
Telescope
Camera
Focusser
Cables for all the above
Pier-cam webcam
Red / white lighting in the dome
White lighting in dome side
Power sockets in dome
Network switch to replace slow hub (for network traffic reasons)
Internal weather sensor junction box
2 internal PIRs
1 external PIR
2 more cloud sensors

The Teide camera was replaced, the internal temperature, humidity and pressure sensors were moved to the new junction box, and various other things were fixed, painted, got rid of, cleaned, etc.

We also started work on aligning the telescope.

Submitted by chris on Friday 30 May 2003

Results from the latest Tenerife trip...

We installed the following: A wind direction sensor, a solar radiation sensor to replace our old light sensor, a bigger and better power supply for the weather station, the optical encoder for the dome, a new ultra-low light camera for pole-star-cam, 2 new UPSs, red lighting, network ports and power sockets.

We fixed the following: the cloud sensors, the dome power supply box, the view from dome-cam, the view from road-cam, the webcams power supply, a window.

Other stuff we did: Tested dome, sorted out UPS software, moved the MOXA unit to it's proper place and recabled it and generally tidied the place and boxed up more old stuff.

Stuff we couldn't do (because IAC people were finishing off our new roof hole): Move internal temperature and humidity sensor to dome side, install mount-cam and install power and lighting in dome.

Overall, a successful trip, except for the dome related jobs.

The next trip out there is in the planning stage.. We hope to set up the first telescope and get it taking images.

Submitted by chris on Thursday 8 May 2003

Web server The new server described in the last post has so far performed perfectly although through my own fault there were some teething problems moving the traffic over from the old server to the new.. ;)

Tenerife We are off to Tenerife again sometime soon to complete all the little jobs that need to be done before we can take out and install the telescopes. This list currently includes things like installing the last part to the dome, installing a wind direction and a better light sensor, replacing the low-light pole star camera with an even lower light camera, moving more weather sensors from old system to new (this includes fixing the cloud sensors), installing some PIR sensors, various wiring and electronical work and a whole host of other miscellaneous jobs. Hopefully the site will be ready to take out the mount and the first telescope in June.

Pole Star Cam Over the last 3 weeks I have developed some replacement software for this camera. It is on a seperate system to the other 3 cameras and now has the ability to sum multiple images, threshold away the background noise, count how many stars it can see and log those results. It is hoped that these results can be used to calibrate the cloud sensors. The new software is also much faster at returning an image to a web browser. It is currently in testing - possibly sometime in the future it will be possible to view the summed and thresholded images over the web.

Submitted by chris on Thursday 17 April 2003

Having argued and waited for months, our new web server arrived from Dell on Monday. As of about half an hour ago it is now running the domain, that is DNS, Web, Email and associated background services. Suffice to say it is a very nice upgrade from the last machine, with it being able to do everything 10 times faster...!
The new machine is a Dell Poweredge 1600SC, 2 2GHz Xeon processors, 2GB RAM, 2 80GB disks in mirror RAID format, running RedHat Linux 8.0.

Submitted by chris on Tuesday 25 March 2003

Still working on buying a new server for the main web site. There is a bit of a lapse from Tenerife work at the moment (!) so I've been filling in various bits of the website code I have always been "leaving till later". The permissions and access rights to various parts of the website are now much better and can support any number of combinations of rights. The site will now correctly respond to what you can and can't do. The security policy on displaying pages has changed - unless all the access rights are set up perfectly the site assumes there is a problem and refuses to serve the page. A definite improvement in security. Next on the list is to rewrite the telescope filter system to be more inline with how most of the rest of the site works, and to implement the teacher/class grouping system for user accounts. Additionally, myself and Dan have begun brainstorming, designing and researching tools for writing the master control program that will manage the telescope site itself. Eek, it ain't simple!

Submitted by chris on Tuesday 18 March 2003

Back from Tenerife again! This time we built the dome, moved some weather sensors to the new weather system minus spikes, started testing some UPS gear, seperated pole star cam from the other cameras so we can use different software with it, and reorganised what the server computers do out there due to the addition of the last server last time we went out there. Looking to the near future for me includes modifying and improving the access rights system for this website, finding and installing different software for pole star cam and possibly ordering a new server for the main website. Again. First though I need to battle through a bit of university red tape to get our Internet access back for my work station and the main web server... (Update:) The Ultra 5 Sun server we use decided to forget it's route since the upstream router went down for a few hours ?!

Submitted by dan on Saturday 22 February 2003

We have just been out to Tenerife, lots of work has been done. I'm not going to fill you in on that just now though, just is just a quick message ot explain the loss of weather data on the morning of Saturdat 22nd. To put it simply the new weather capture software (which is very very good) wasn't expecting an ntp client to run at 5.15 in the morning. When the time client ran it changed the system clock. The weather station does samples based on a very accurate 10 second timer. When the ntp software ran it changed the system clock, so when the next sample was taken the time was not within the 1 second tolerance to allow addition to the database. So since the ntp client changed the clock, no data has entered the database. It's a quite simple fix, but it'll wait till monday morning. (We only arrived back in the UK at about 4:15, and with the speedy work of the luggage handlers at Manchester it only took and hour and thirty minutes to get out of the airport)

Submitted by chris on Thursday 30 January 2003

Things have been moving fairly slowly recently, but we are now preparing for the next trip out to Tenerife. We have ordered all the parts to build some new weather station boxes to replace parts of the old weather system out there. We have been tracking down the fault with the weather data - it contains large negative spikes at strange intervals. We will be attempting to fix that. Other things we are preparing include a video switch box to allow many cameras to be fed into one capture card to simplify the webcam system. We have also purchased a Moxa serial server to enable us to have the multitude of serial devices we plan to use in the final system. The primary reason for this trip to Tenerife, however, is to install the metal work required to support the new dome roof.

Submitted by chris on Monday 16 December 2002

Having finished our sub-project (non-telescope related) we have returned our attention to BRT business. We spent a week preparing for a trip to Tenerife and last week we went out there. What we did: Replaced the Road-Cam camera with a new hopefully sun-resistant camera. Installed Teide-Cam. Half installed Pole-star-Cam. Rewired the camera power system to support several cameras. Purchased, installed and installed software on a new computer - the computer to run TheSky and talk to the telescope mount. Installed conduiting for cable runs around the building. The other team of people dismanted the old roof opening system and covered the resulting hole, ready for the new dome roof to be installed. They also sealed various points of the building that might or were already leaking.

Submitted by chris on Wednesday 20 November 2002

Over the last month I have been more involved with non-telescope.org work, however in my spare time (!) I have rewritten the site protocol software. This is the software that sits on the main webserver that talks to the actual telescope sites, and the software at the site that talks back. The first implementation worked, but was not easy to maintain and update. It has provided vital information on how the system should work though and was a useful stepping stone to the current software which is a significant improvement.
For those interested, the old system was a daemon written in C on site and php client scripts. The new software is a PHP daemon on site instead. This makes it much easier to maintain and update the software, it makes it easier to guarantee that the code is safe and has some very useful and easy-to-use data encapsulation and database modules. There are now enough modules in PHP to create a real(ish) daemon, and the fact that it all runs interpreted is irrelevant as the processing capacity to support it is present. (On the site side at least).

Submitted by chris on Friday 18 October 2002

Over the last few weeks I have mostly been working on all the things that I left until later to complete - either small things that didn't need to be done at the time or more difficult things that I tried to ignore! Some examples... The SAO, NGC and IC catalogues are now in the system and are available for selection, the main server automatically backs up all important data to another machine periodically and I wrote a page hit system which automatically works on every page. I looked into why the weather page took so long to load and with some code changes I have halved the time it takes to do so (it could still do with being faster though). I finally got around to changing all code that deals with user submitted data to use the PHP super globals arrays rather than using register_globals. This cleared up a few known security problems and probably eliminated many more future ones. The site is now immune to people submitting the same form twice. It turns out this is only a problem on pages that insert new rows into tables - pages that modify rows or delete rows will not corrupt any data if the form is resubmitted. However, if a page that inserted data into a table was run twice then it was possible to generate the same row in the database. This didn't corrupt the database because the rows receive unique ID numbers but it did corrupt the logic of the program. This is fixed by handing out a random number ticket to the browser when it displays the form. If that ticket number does not come back with the submitted data then the data is ignored. Once that ticket is returned it is invalidated and cannot be used again. I have written some patch code into pages affected by an oddity in IE where sometimes the value of a button is returned and sometimes it isn't. And lastly I am finally writing in code wherever objects are deleted to make sure no other data is affected or invalidated. These items along with many other smaller tasks are making the site far more bullet proof.

Submitted by dan on Friday 11 October 2002

Here comes the sun. Has anyone noticed anything wrong with the web cam? Big streaky lines in the top left. Have you guessed it yet? Well the sun is now passing across the field of view of the camera; it is so intense that it is damaging the pixels of the CCD within the camera. Fear not though, we have a plan in testing to fix this problem.

Later on this month there is a trip planned to the site in Tenerife, the primary role of this trip is to make the site ready for the installation of our new dome at the start of next year. In addition to this work there will also be two new web cameras added, as well as repairs made tot he current one.
The first of the new web cams will hopefully be an ultra low light monochrome camera, which will be pointed at the pole star. The second of the new cameras will be a standard colour camera like the existing one, but this time pointed at mt. Teide.

In addition to all of the work being planned for Tenerife there has also been progress made on the software control systems for the remote sites. The main advance of the new software is a modular structure with dynamic loading capabilities. To the user this will be invisible, but will allow us to update hardware and software systems without the site going offline. This will in part allow us to give a higher quality of service with less downtime. In addition it should give us a higher capacity and help to minimise the waiting period for observations.

Submitted by chris on Friday 20 September 2002

It's taken until this week to get the server situation truly sorted out. I have had to stay with Sun hardware running Solaris rather than ordering a new server that I would have run Linux on. So I tried to replace the faulty disk in the Ultra 10 but it didn't work due to incompatibilities we havn't tracked down yet. All attempts to run the Ultra 10 with a different hard disk failed for some reason or another, so I was given two Ultra 5 computers instead. They are newer and each slightly more powerful than the Ultra 10 was so I thought I would put DNS, Sendmail and Postgres on one and Apache/PHP on the other. However one of them crashed while compiling software and crashed again over last weekend when it was sat there doing nothing, so I've stacked that with the broken Ultra 10... So now we have one Ultra 5 running everything. We'll see how it goes. In lighter news I swapped the website to this new version on Wednesday which is a big mental step forward in the project for me. Since Wednesday I have been ironing out small problems with the site move, getting the weather pages working properly and doing various other small server organising tasks. There are still a few holes in the web site software which I will work on over the next few weeks in parallel with writing up my MPhil. The site will basically now wait until the telescope hardware side of things catches up.

Submitted by chris on Friday 30 August 2002

Bad weather spotted at the Tenerife observatory site! A rare thing during summer... Webcam image from the time follows, as well as snapshots of the Bradford Robotic Telescope weather data at the time and the GONG project weather data at the time. They are gratifyingly similar. The snapshots will need Internet Explorer 4, or 5 perhaps.



Bradford Robotic Telescope Weather Data snapshot

GONG project weather data snapshot

Submitted by chris on Wednesday 31 July 2002

Yet again, far too much to remember but here's some of it.
I have changed the forum software to phpBB. It can now use Postgres as a backend database, which is what the rest of the system uses.
I have been working on the job submission scripts, improving / rewriting and improving again. One day they will do everything they are supposed to and will be finished...
Other than that I have been filling in bits of scripts that can now be written.
Last week we went to Tenerife again (read Dan's post below).
On returning from Tenerife I discovered that one of the disks in the main telescope.org server doesn't start up well after power outages. I coaxed it into working again but I think it's time to get a big powerful machine to handle the website (as I think it will need something more than what we have now when the system actually starts working.) So I will be ordering and installing that soon, and then it's time to put this website on www instead of dev!

Submitted by dan on Monday 29 July 2002

This is a really really long news update... so far..
* We now have some server software under windows to move the telescope to objects and locations.

* We have some software under windows to slew the telescope to objects and locations using the server.

* A c++ class have been started to talk to the server from linux.

* New humidity sensor arrived from Skye Instruments. (It's the correct sensor too... which is nice.. there was some worry they wouldn't be able to make a specially calibrated version of the sensor in time.)

* We went out and made a nice new toolkit for the trip to tenerife.

* We went to tenerife... and did all this...

--* Discovered that our car had no radio or air con.

--* Replaced the humidity sensor.

--* Moved and tuned the cloud sensors.

--* Tested the cloud sensor circuits.

--* Removed most of the old telescope wiring.

--* Cleaned up the building.

--* Purchased conduit to clean up cabling in new installation.

It think thats more/less all... apart from the fact I miss the laptop :(

Submitted by chris on Monday 1 July 2002

Wow. Long time since last update! Too long to remember everything we've done! But here is some of it. We have made several trips to the Oxenhope site to get the ISDN working (again) and to spend time cleaning out the hut. Since the ISDN is so unreliable we have installed a landline dial-in service so that we can access the site even if the ISDN is down. While testing and wiring phone lines we appear to have improved the signal quality on the main site line to a level capable of about a 20kbps modem connection! On the latest trip we installed a new weather station.
We have ressurected a dead ups we brought back from Oxenhope, so that can be used to allow the Meade LX-200 to find its home position in case of a mains failure.
Myself and John went to Tenerife to do various jobs. We cleared out half of the telescope container and installed one new PC. We installed a web cam and reconnected all the old weather station (still working!). The web cam picture and the weather data can be seen on this website.
Baldrick has been moved to the lab from a server room and is no longer providing any telescope.org services.
I reorganised our IP addresses and DNS names due to the ever growing number of machines we have.
And finally, I went on holiday for a week. It was nice.

Submitted by chris on Friday 3 May 2002

Forums It was taking too long to find perfectly suitable forum software so I have installed YapBB. The whole forums.telescope.org site runs on a different server so as not to taint the main server with MySQL :)
Web Site Move Since Baldrick the old server has to move out of the room it is in, I have had to move forward the plan to move the website over to the new box (Gandalf). I have done a reasonably faithful mirror of all the non-dynamically generated content onto the new box and have changed the DNS records to send traffic going to www.telescope.org to Gandalf instead of Baldrick. As of today most of the traffic is now hitting the new box.
Control Protocol I finally started to look into this and researched and considered XML for the protocol but this seemed like a non-viable idea so I am looking into other ways of doing it. From this I sidetracked off onto some database issues that really needed to be solved first... Which leads me onto
Postgres and PHP Upgrade Both the development box and the main box now have Postgres 7.2.1 and PHP 4.2.0 so that I can play with the new bytea type for storing data.

Submitted by chris on Friday 19 April 2002

Time for my two-weekly update!

Web Site
I've finished the tracker so we can record, track and keep in touch with what we're trying to do. Once the tracker was finished Iain the beta-tester filled it with quite a few bugs, which I set about fixing, and have fixed most of now. I have written the mentioned blogger, (the system to allow us to dynamically update this news page).

Database
I spent quite a lot of time rewriting code to a more efficient and flexible design. This could only be done now as I could not see before how parts of the code neede to operate, whereas I do now. With PHP not understanding Postgres' boolean type properly I also had to redesign some code to use standard numbers instead...

Forums
Difficult one. I run Postgres for everything, so I don't want to run MySQL just for a webboard. I looked into forums and of course the best ones need MySQL. I looked at converting one to use Postgres, and while most of it was easy there were one or two killer problems, making it nearly impossible. Still working on it though, currently I'm trying to obtain one that looks not too bad and runs with Postgres.

Servers, networking
Both the main server and the development server are now correctly time syncing, which is nice. I obtained a subnet so we can route IP traffic to the Oxenhope site. And finally I am currently working on retrieving the old sprawling website and putting it on a new server so that Baldrick the old server can be shut down.

Submitted by dan on Thursday 18 April 2002

On the 15th of April three new computers were ordered for the testing and development of the on site elements of the telescope control system. Hopefully two of these will run linux, operating the main control system, database, focusers, filters and ccd imagers. The remaining machine will be running Windows of some NT variety to acutally control the telescope mount (its a win-mount... oh well).

On the other side of things, yay, hopefully not long now till the mount control software arrvies. Due to a quite cool stroke of luck the software that controls the new Paramount ME mounts also can control the Meade LX200 which is currently in the lab. As soon as the software arrives we can connect it up to the LX200 and it will provide and identical control interface to the one that will be used with the new mounts. So the long and the short of it is that we can start developing the software for the control of the mount even though it wont arrive for another few months.

After all this good news its time to break the new so good news. To be blunt the LX200 is stupid. If it is powered down then powered up it assumes that its pointing in its park direction. In other words it has no hardware reference point to align itself from. This means in the event of a power failure the telescope cannot safely bring itself back online without intervention on site. I am currently looking into ways of connecting the telescope control computer and telescope mount to a UPS to allow it to home before power failure to solve this little problem.

Submitted by dan on Friday 12 April 2002

Contacted Software Bisque at about 5:30pm on the 11th, the order has been recieved and is ok. yippee. The waiting time for the mounts is around the six month mark. On the brighter side they are currently having new CDs of the software made, and the CDs should arrive in one to two weeks. Once the software has arrived we can start more serious development of the control software for the mount using the Meade LX200.

Submitted by chris on Friday 5 April 2002

Ok, so no entries for ages, even though things have happened. But hey, nobody was reading this page because a) you couldn't get to it easily from the outside of the university, and b) nobody knew the url anyway. (I'll try to update this more often now.) Anyhoo.. As of Monday 18 March both the development web server and the main base station server's web server can be accessed from the outside, and with good urls. The development server is dev.telescope.org and the main server is new.telescope.org - which will change to www.telescope.org when I can finally swap over the servers. What has happened since the last update? DNS is done, and the hosting service has been moved to the new server. The smtp mail service has been moved over to the new server. We've got a big lab now to put telescopes and weather stations in, and we received a Meade LX-200 today to play with. Things do seem to be rolling. A bug tracking system is nearly finished, which should help us track ideas and bugs. I might set up a web board thing so we can have open public rolling discussions with people, or I might not. It might not be politically correct either to do that on a university server. Also, this page looks really crowded and silly now, I might actually need some kind of blog system (wow!).

Submitted by chris on Wednesday 6 February 2002

DNS is no closer. You'd think it would be simple to change the two DNS servers listed at Network Solutions that host the domain, but no. I have been battling with them for weeks now and the only thing I have accomplished is that I am technical contact for the domain. Which does me no good because I can't assign the two new DNS servers to the domain because they are either not listed or have the wrong IP address. Which I need other people to solve. The battle goes on.

Submitted by chris on Monday 28 January 2002

New Sun UltraSPARC server finished being installed. Placed, plugged in and ready to go. The fight goes on to get Network Solutions to change the DNS servers for the domain.

Submitted by chris on Thursday 20 December 2001

Development of new web site began 12 December 2001.

I hope to have the mail and DNS system on telescope.org revamped soon.

University of Bradford
PayPal
Help us by donating
About us
Contact us
Site map
Statistics
Photo galleries
Forum
System status
Telescope specs
My Account
FAQs
© 2015 University of Bradford. Terms of use, privacy policy